how do I solve 10x+11=-19
Answer by unlockmath(1688)

Hello,
Whatever you do on one side of the equal sign you have to do the exact same thing on the other side.
OK, the trick is to get x all alone on one side of the equal sign.
So first, let's move +11 to the other side by subtracting 11

10x + 11 - 11 = -19 - 11
That leaves us with:
10x = -30
Now, divide each side by 10
10x/10 = -30/10 (remember, you have to do the
same on both sides)
That gives us
x = -3
You can always check the answer by plugging it into the original equation.
RJ Toftness
Author of "Unlock the Mystery to Math"
www.math-unlock.com